MusicGeek:

i need them all answered except for 17 and 18

MusicGeek:

Narad:

Area of circle \[A=\pi R^{2}\] Area of sector of a circle \[a=\frac{ 1 }{ 2 }\pi R^{2}\] Circumference of a circle \[=2 \pi R\] Length of arc of a circle \[=R \theta \]

MusicGeek:

what is θ and what does it mean and does it have any numerical value?

Narad:

\[\theta \] is the angle PRQ

Narad:

Angle subtended by the arc PQ